# Predicting the Performance of Black-box LLMs through Self-Queries

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance on a wide variety of tasks (Radford et al., ), leading to their increased involvement in larger systems. For instance, they are often used to provide supervision (Bai et al., 2022; Sam et al., 2024), as tools in decision-making (Benary et al., 2023; Sha et al., 2023), or as controllers on agentic frameworks (Xi et al., 2023; Robey et al., 2024). Thus, it is crucial to understand and predict their behaviors, especially in high-stakes settings. However, as with any deep network, it is difficult to understand the behavior of such large models (Zhang et al., 2021). For instance, prior work has studied input gradients or saliency maps (Simonyan et al., 2013; Zeiler & Fergus, 2014; Pukdee et al., 2024)) to attempt to understand neural network behavior, but this can fail to reliably describe model behavior (Adebayo et al., 2018; Kindermans et al., 2019; Srinivas & Fleuret, 2020). Other work has studied the ability of transformers to represent certain algorithms (Nanda et al., 2022; Zhong et al., 2024) that may be involved in their predictions.
@@ 3.2 Generating Follow-up Prompts
We note that based on the specific nature of the question, the response (e.g., the probability of responding yes) could define a weak predictor of whether the model is correct or not. This is reminiscent of the design of weak learners in boosting (Freund & Schapire, 1996) or weak labelers in programmatic weak supervision (Ratner et al., 2017; Sam & Kolter, 2023; Smith et al., 2024). However, to maintain our approach’s generality and to not restrict our approach to only a certain type of elicitation questions, we treat these as abstract features for a linear predictor. We also note that further work could perform discrete optimization over prompts to further improve the extracted representation’s usability, through methods described in (Wen et al., 2024; Zou et al., 2023b; Chao et al., 2023). However, one key appeal of the current approach is that it defines an extremely simple classifier in a task-agnostic fashion. Performing optimization over these questions might lead to overfitting, and the resulting predictors on the outputs of these prompts require more complex analysis in deriving valid generalization bounds.
@@ 3.3 Analysis on Finite Samples from Black-box LLMs
Recall that we have our representation $z=(z_{1},...,z_{d})$, which corresponds to the actual probability of the yes token under the LLM. Without access to these true probabilities through an API, we instead have some approximation $\hat{z}=(\hat{z}_{1},...,\hat{z}_{d})$, where each $\hat{z}_{j}$ is an average of $k$ samples from $\text{Ber}(z_{j})$. From prior work in logistic regression under settings of covariate measurement error (Stefanski & Carroll, 1985), when we have that $k$ grows with $n$, we observe that the naive MLE (maximum likelihood estimator) on the observed approximation results in a consistent, albeit biased, estimator. We present an analysis of our setting, showing a result on the convergence rate of the MLE for $\beta$.
@@
Figure 2: AUROC in predicting model performance on the open-ended QA benchmarks of Natural Questions (Top) and SQuAD (Bottom). Dashed bars represent black-box methods, which assume more access than QueRE. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for the GPT models is an approximation of a sparse vector with nonzero values for the top-5 logits from the API.
Figure 3: AUROC in predicting model performance on closed-ended QA benchmarks of HaluEval, BoolQ, and DHate. Dashed bars represent black-box methods. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for GPT models is a sparse vector with nonzero values for the top-5 logits from the API.
@@ 4.4 Additional Results
Another added benefit of our approach is that it yields low-dimensional representations, which can be used with simple models, to achieve strong predictors of performance with tight generalization bounds. Bounds for linear models that use features from a pretrained model have been explored in practice (McNamara & Balcan, 2017), although not for LLMs. Another key difference is that, while we similarly extract a representation from the model, previous approaches use a penultimate layer rather than the ability of a LLM to generated features in response to language queries. We use the following PAC-Bayes generalization bound for linear models (see Section A.5 for more details). We observe that linear predictors trained our representations have stronger guarantees on accuracy, when compared to baselines (Table 3 and Appendix A.5). A limitation of these results is that they require an assumption that the representations extracted by a LLM are independent of the downstream task data; this assumption is verifiable via works in data contamination (Oren et al., 2023) or is valid on datasets released after LLM training (e.g., HaluEval for GPT-3.5).
@@ A.2 Uncertainty Quantification Baselines
Another line of work in uncertainty quantification (Xiong et al., 2023) looks to extract estimates of model confidence from the LLM directly. This is fundamentally related to our problem setting, but perhaps is less focused on the applications of predicting model behavior (and certainly not focused on our other applications of detecting adversarial models or distinguishing between architectures). These baselines include: (1) Vanilla confidence elicitation, which is to directly ask the model for a confidence score, (2) TopK, asking the LLM for its TopK answer options with their corresponding confidences, (3) CoT, asking the LLM to first explain its reasoning step-by-step before asking for a confidence score, and (4) Multistep, which asks the LLM to produce multiple steps of reasoning each with a confidence score. We use $K=3$ for the TopK baseline and 3 steps in the multistep baseline.
@@
We observe that QueRE achieves stronger performance than these these uncertainty quantification baselines (Table 6). We also remark that QueRE is more widely applicable as these methods (which are implemented in Xiong et al. (2023)), as they heavily on being able to parse the format of responses for closed-ended QA tasks. On the contrary, QueRE indeed applies to open-ended QA tasks (see our strong results in Figure 2).
@@ A.5 Additional Generalization Results
For our PAC-Bayes bounds over linear models (Jiang et al., 2019), we use a prior over weights of $\mathcal{N}(0,\sigma^{2}I)$, giving us our bound as
@@
We remark that our work defines a different line to approach generalization bounds through a more human-interactive approach to eliciting low-dimensional representations, although we remark that this human-interaction in specifying these elicitation questions must be independent of any training data (e.g., questions must be predefined before seeing the dataset of interested). Perhaps the most related work in this line are existing works that have achieved tight generalization bounds for VLMs (Akinwande et al., 2023) and for LLMs modeling log-likelihoods (Lotfi et al., 2023).
@@ A.10 Additional Results for Varying the Number of Elicitation Questions
We present additional results when varying the number of elicitation questions on other QA tasks. Here, we only look at subsets of the elicitation questions and do not include the components of preconf, postconf and answer probabilities. We observe that across all tasks, we observe a consistent increase in performance as we increase the size of the subset of elicitation questions that we consider, with diminishing benefits as we have a larger number of prompts (Figure 13). Generally, increasing the number of elicitation prompts leads to an increase in AUROC, clearly defining a tradeoff between extracting the most informative black-box representation and the overall cost of introducing more queries to the LLM API. An interesting future question is how to best select elicitation questions, and perhaps, removing those that add redundant information or noise. This is reminiscent of work in prior work in pruning or weighting ensembles of weak learners (Mazzetto et al., 2021a; b) or in dimensionality reduction (Van Der Maaten et al., 2009).
@@ Appendix B Proof of Proposition 1
Consider the standard logistic regression setup (as in the work of Stefanski & Carroll (1985)), where we are learning a linear model $\beta$, which satisfies that
@@
In our setting, instead of having access to covariates $X_{i}$, we rather have access to an approximation of these covariates $\hat{X}_{i}$, which is an average of $k$ samples from Ber($X_{i}$). An application of the results in the work of Stefanski & Carroll (1985) gives us the result that the MLE $\hat{\beta}$ is a consistent estimator of $\beta_{0}$, given that $k\to\infty$. This is fairly straightforward as when $k\to\infty$, we have that $\frac{1}{k}\sum_{j=1}^{k}\hat{X}_{i}^{j}\to X_{i}$, implying that the noise in the covariates goes to 0 as $n\to\infty$ (i.e., satisfying a main condition of the result in Stefanski & Carroll (1985)).
@@ Appendix C Additional Related Work
A large body of work has focused on understanding the capabilities of LLMs. The field of mechanistic interpretability has recently evolved around understanding the inner workings of LLMs by uncovering circuits or specific weight activations (Olsson et al., 2022; Nanda et al., 2022). This has developed a variety of potential hypotheses for how models learn to perform specific tasks (Zhong et al., 2024), as well as the tendencies of certain activations in a LLM to activate on certain types of inputs (Bills et al., 2023; Sun et al., 2024). Other works have studied model behavior by locating specific regions of a LLM that relate to certain concepts such as untruthfulness (Campbell et al., 2023) or honesty and ethical behavior (Zou et al., 2023a). Our work is different in that we only assume black-box access, with a similar goal to extract information about model behavior. Finally, other work has attempted to study the abilities and performance of LLM via developing challenging benchmarks (Hendrycks et al., 2020), also including those that use techniques from the cognitive sciences (Binz & Schulz, 2023) or by comparing with human similarity judgments (Coda-Forno et al., 2024). While these approaches look to benchmark and quantify performance in aggregate over tasks, our setting looks to predict the performance at the example level, for deciding when to trust or use LLMs in deployment.
@@ D.2 Adversarial Code Experiment Details
We now present additional details for our adversarially influenced code generation experiment in Section 4.2. In our setting, we take 500 train examples of coding prompts from the Code Contests dataset (Li et al., 2022) and evaluate on their full test set. We evaluate the ability of linear predictors trained on top of QueRE and alternative approaches to detect adversarially influenced models that behave normally when answering normal questions but introduce bugs into generated code.
@@ D.7 Generalization Details
For our generalization details, we use PAC-Bayesian bounds over the linear models, as is outlined in the work of Jiang et al. (2019). Here, we consider a prior of weights specified about the origin, with a grid of variances of [0.1, 0.11, …, 0.99, 1.0]. For the generalization experiments, we balance both the train and test datasets as we evaluate the accuracy of different predictors.

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance on a wide variety of tasks (Radford et al., 2019), leading to their increased involvement in larger systems. For instance, they are often used to provide supervision (Bai et al., 2022; Sam et al., 2024), as tools in decision-making (Benary et al., 2023; Sha et al., 2023), or as controllers on agentic frameworks (Xi et al., 2023; Robey et al., 2024). Thus, it is crucial to understand and predict their behaviors, especially in high-stakes settings. However, as with any deep network, it is difficult to understand the behavior of such large models (Zhang et al., 2021). For instance, prior work has studied input gradients or saliency maps (Simonyan et al., 2013; Zeiler and Fergus, 2014; Pukdee et al., 2024)) to attempt to understand neural network behavior, but this can fail to reliably describe model behavior (Adebayo et al., 2018; Kindermans et al., 2019; Srinivas and Fleuret, 2020). Other work has studied the ability of transformers to represent certain algorithms (Nanda et al., 2022; Zhong et al., 2024) that may be involved in their predictions.
@@ 3.2 Generating Follow-up Prompts
We note that based on the specific nature of the question, the response (e.g., the probability of responding yes) could define a weak predictor of whether the model is correct or not. This is reminiscent of the design of weak learners in boosting (Freund and Schapire, 1996) or weak labelers in programmatic weak supervision (Ratner et al., 2017; Sam and Kolter, 2023; Smith et al., 2024). However, to maintain our approach’s generality and to not restrict our approach to only a certain type of elicitation questions, we treat these as abstract features for a linear predictor. We also note that further work could perform discrete optimization over prompts to further improve the extracted representation’s usability, through methods described in (Wen et al., 2024; Zou et al., 2023b; Chao et al., 2023). However, one key appeal of the current approach is that it defines an extremely simple classifier in a task-agnostic fashion. Performing optimization over these questions might lead to overfitting, and the resulting predictors on the outputs of these prompts require more complex analysis in deriving valid generalization bounds.
@@ 3.3 Analysis on Finite Samples from Black-box LLMs
Recall that we have our representation $z=(z_{1},...,z_{d})$, which corresponds to the actual probability of the yes token under the LLM. Without access to these true probabilities through an API, we instead have some approximation $\hat{z}=(\hat{z}_{1},...,\hat{z}_{d})$, where each $\hat{z}_{j}$ is an average of $k$ samples from $\text{Ber}(z_{j})$. From prior work in logistic regression under settings of covariate measurement error (Stefanski and Carroll, 1985), when we have that $k$ grows with $n$, we observe that the naive MLE (maximum likelihood estimator) on the observed approximation results in a consistent, albeit biased, estimator. We present an analysis of our setting, showing a result on the convergence rate of the MLE for $\beta$.
@@
Figure 2: AUROC in predicting model performance on the open-ended QA benchmarks of Natural Questions (Top) and SQuAD (Bottom). Dashed bars represent white-box methods, which assume more access than QueRE. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for the GPT models is an approximation of a sparse vector with nonzero values for the top-5 logits from the API.
Figure 3: AUROC in predicting model performance on closed-ended QA benchmarks of HaluEval, BoolQ, and DHate. Dashed bars represent white-box methods. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for GPT models is a sparse vector with nonzero values for the top-5 logits from the API.
@@ 4.4 Additional Results
Another added benefit of our approach is that it yields low-dimensional representations, which can be used with simple models, to achieve strong predictors of performance with tight generalization bounds. Bounds for linear models that use features from a pretrained model have been explored in practice (McNamara and Balcan, 2017), although not for LLMs. Another key difference is that, while we similarly extract a representation from the model, previous approaches use a penultimate layer rather than the ability of a LLM to generated features in response to language queries. We use the following PAC-Bayes generalization bound for linear models (see Section A.5 for more details). We observe that linear predictors trained our representations have stronger guarantees on accuracy, when compared to baselines (Table 3 and Appendix A.5). A limitation of these results is that they require an assumption that the representations extracted by a LLM are independent of the downstream task data; this assumption is verifiable via works in data contamination (Oren et al., 2023) or is valid on datasets released after LLM training (e.g., HaluEval for GPT-3.5).
@@ A.2 Uncertainty Quantification Baselines
Another line of work in uncertainty quantification [Xiong et al., 2023] looks to extract estimates of model confidence from the LLM directly. This is fundamentally related to our problem setting, but perhaps is less focused on the applications of predicting model behavior (and certainly not focused on our other applications of detecting adversarial models or distinguishing between architectures). These baselines include: (1) Vanilla confidence elicitation, which is to directly ask the model for a confidence score, (2) TopK, asking the LLM for its TopK answer options with their corresponding confidences, (3) CoT, asking the LLM to first explain its reasoning step-by-step before asking for a confidence score, and (4) Multistep, which asks the LLM to produce multiple steps of reasoning each with a confidence score. We use $K=3$ for the TopK baseline and 3 steps in the multistep baseline.
@@
We observe that QueRE achieves stronger performance than these these uncertainty quantification baselines (Table 6). We also remark that QueRE is more widely applicable as these methods (which are implemented in Xiong et al. [2023]), as they heavily on being able to parse the format of responses for closed-ended QA tasks. On the contrary, QueRE indeed applies to open-ended QA tasks (see our strong results in Figure 2).
@@ A.5 Additional Generalization Results
For our PAC-Bayes bounds over linear models [Jiang et al., 2019], we use a prior over weights of $\mathcal{N}(0,\sigma^{2}I)$, giving us our bound as
@@
We remark that our work defines a different line to approach generalization bounds through a more human-interactive approach to eliciting low-dimensional representations, although we remark that this human-interaction in specifying these elicitation questions must be independent of any training data (e.g., questions must be predefined before seeing the dataset of interested). Perhaps the most related work in this line are existing works that have achieved tight generalization bounds for VLMs [Akinwande et al., 2023] and for LLMs modeling log-likelihoods [Lotfi et al., 2023].
@@ A.10 Additional Results for Varying the Number of Elicitation Questions
We present additional results when varying the number of elicitation questions on other QA tasks. Here, we only look at subsets of the elicitation questions and do not include the components of preconf, postconf and answer probabilities. We observe that across all tasks, we observe a consistent increase in performance as we increase the size of the subset of elicitation questions that we consider, with diminishing benefits as we have a larger number of prompts (Figure 13). Generally, increasing the number of elicitation prompts leads to an increase in AUROC, clearly defining a tradeoff between extracting the most informative black-box representation and the overall cost of introducing more queries to the LLM API. An interesting future question is how to best select elicitation questions, and perhaps, removing those that add redundant information or noise. This is reminiscent of work in prior work in pruning or weighting ensembles of weak learners [Mazzetto et al., 2021a, b] or in dimensionality reduction [Van Der Maaten et al., 2009].
@@ Appendix B Proof of Proposition 1
Consider the standard logistic regression setup (as in the work of Stefanski and Carroll [1985]), where we are learning a linear model $\beta$, which satisfies that
@@
In our setting, instead of having access to covariates $X_{i}$, we rather have access to an approximation of these covariates $\hat{X}_{i}$, which is an average of $k$ samples from Ber($X_{i}$). An application of the results in the work of Stefanski and Carroll [1985] gives us the result that the MLE $\hat{\beta}$ is a consistent estimator of $\beta_{0}$, given that $k\to\infty$. This is fairly straightforward as when $k\to\infty$, we have that $\frac{1}{k}\sum_{j=1}^{k}\hat{X}_{i}^{j}\to X_{i}$, implying that the noise in the covariates goes to 0 as $n\to\infty$ (i.e., satisfying a main condition of the result in Stefanski and Carroll [1985]).
@@ Appendix C Additional Related Work
A large body of work has focused on understanding the capabilities of LLMs. The field of mechanistic interpretability has recently evolved around understanding the inner workings of LLMs by uncovering circuits or specific weight activations [Olsson et al., 2022, Nanda et al., 2022]. This has developed a variety of potential hypotheses for how models learn to perform specific tasks [Zhong et al., 2024], as well as the tendencies of certain activations in a LLM to activate on certain types of inputs [Bills et al., 2023, Sun et al., 2024]. Other works have studied model behavior by locating specific regions of a LLM that relate to certain concepts such as untruthfulness [Campbell et al., 2023] or honesty and ethical behavior [Zou et al., 2023a]. Our work is different in that we only assume black-box access, with a similar goal to extract information about model behavior. Finally, other work has attempted to study the abilities and performance of LLM via developing challenging benchmarks [Hendrycks et al., 2020], also including those that use techniques from the cognitive sciences [Binz and Schulz, 2023] or by comparing with human similarity judgments [Coda-Forno et al., 2024]. While these approaches look to benchmark and quantify performance in aggregate over tasks, our setting looks to predict the performance at the example level, for deciding when to trust or use LLMs in deployment.
@@ D.2 Adversarial Code Experiment Details
We now present additional details for our adversarially influenced code generation experiment in Section 4.2. In our setting, we take 500 train examples of coding prompts from the Code Contests dataset [Li et al., 2022] and evaluate on their full test set. We evaluate the ability of linear predictors trained on top of QueRE and alternative approaches to detect adversarially influenced models that behave normally when answering normal questions but introduce bugs into generated code.
@@ D.7 Generalization Details
For our generalization details, we use PAC-Bayesian bounds over the linear models, as is outlined in the work of Jiang et al. [2019]. Here, we consider a prior of weights specified about the origin, with a grid of variances of [0.1, 0.11, …, 0.99, 1.0]. For the generalization experiments, we balance both the train and test datasets as we evaluate the accuracy of different predictors.